# Shamier Settle, Matthew Shapiro, Emerita Torres, and Anand Vidyashankar for discussion and feedback on the work described

### v5

@@ Abstract
We estimate the number of street vendors in New York City. We first summarize the process by which vendors receive licenses and permits to legally operate in New York City. We then describe a survey that was administered by the Street Vendor Project while distributing Coronavirus relief aid to vendors operating in New York City both with and without a license or permit. Finally, we calculate the total number of vendors using ratio estimation. We find that approximately 23,000 street vendors operate in New York City: 20,500 mobile food vendors and 2,300 general merchandise vendors. One third are located in just six ZIP Codes: 11368 (16%), 11372 (3%), and 11354 (3%) in North and West Queens and 10036 (5%), 10019 (4%), and 10001 (3%) in the Chelsea and Clinton neighborhoods of Manhattan. We also provide a theoretical justification of our estimates based on the theory of point processes and a discussion of their accuracy and implications. In particular, our estimates suggest the American Community Survey fails to cover the majority of New York City street vendors.
@@ 1. Introduction
Street vendors are New York City’s smallest businesses, selling food and merchandise from carts, stalls, and trucks throughout the five boroughs. They are an iconic part of the urban landscape and a thriving sector of the local economy, contributing millions of dollars in government revenue through taxes, fines, and fees. Perhaps most importantly, street vending historically benefits underserved communities, both because vendors operate in neighborhoods with limited access to traditional stores and because vending is one of a handful of occupations in which New Yorkers of all backgrounds, immigrants in particular, can achieve economic mobility and a chance at the American dream (Burrows and Wallace 1998, chap. 42).
@@
In Section 3, we use the fact that the number of licenses and permits are limited by law to construct a ratio estimator for the number of street vendors. We then develop a theoretical justification based on the theory of point processes, which assumes the spatial distribution of survey respondents is representative and well-approximated by a set of inhomogeneous Poisson processes. Finally, we derive the standard errors.
In Section 4, we examine the sensitivity of the ratio estimator to deviations from the underlying assumptions. We first consider the case in which the spatial distribution of the survey respondents is not representative, and we examine a more general estimator in which sampling weights adjust for differential nonresponse. We then consider the case in which the Poisson process assumption is violated, and we examine a more general model in which the arrival of street vendors follows a pure birth process, of which the Poisson and Yule processes are special cases. The Yule process may be more appropriate if street vendors cluster within markets. In both cases, we rederive the standard errors.
In Section 5, we estimate the number of street vendors in New York City. We find approximately 23,000 street vendors operate in New York City: 20,500 mobile food vendors and 2,400 general merchandise vendors. We also find that this estimate is somewhat insensitive to deviations from our underlying assumptions. For example, we find that if the spatial distribution of survey respondents is not representative, the estimated number of street vendors in New York City could be as low as 21,000 and as high as 32,000.
In Section 6, we discuss some implications of our estimates. In particular, we find that although the spatial distribution of respondents matches the results of the American Community Survey, the two disagree on the size of New York City’s street vending population. Our estimates suggest the American Community Survey misses the majority of New York City street vendors. This finding is important because the American Community Survey is commonly used for policy and advocacy, and our estimates shed light on the broader challenges of measuring the “gig economy.”
2. Background and Survey Design
A street vendor is any individual who sells goods from a mobile vending unit instead of a store. We distinguish between the mobile vending unit from which the goods are sold (i.e., the vending establishment) and the individuals who own and/or operate that unit (i.e., the vendors). Note that an establishment refers to a single vending unit. There may be multiple vendors associated with any establishment, and multiple establishments may be owned by a single individual or firm. For example, suppose Sally owns the business Sally’s Salads, which consists of two food trucks, each staffed by three employees. Then Sally’s Salads counts as one business, two establishments, and seven vendors.
@@ 2.1 Background
A long list of rules determines the individuals, locations, and time periods during which vendors can legally sell their goods in New York City. Most relevant is the fact that a merchandise vendor requires a license to operate legally, and a food vendor requires both a license and a permit. The number of food permits and merchandise licenses are limited by law: 5,100 permits and 853 general merchandise licenses are available. Mobile food vending licenses are not limited. Note that a mobile food vending permit is issued to an individual or business to allow for the sale of food from a specific mobile food vending unit, such as a cart or truck. A mobile food vending license authorizes an individual to prepare or serve food from a permitted mobile food vending unit.
There are several versions of a mobile food vendor permit with varying restrictions. Of the 5,100 permits available, 200 are borough permits that limit vendors to one of the five boroughs; 100 are reserved for veterans or vendors with a disability; 1,000 are seasonal and valid only from April to October; 1,000 are green cart permits that limit vendors to selling fruit, vegetables, plain nuts, and water; and 2,800 are unrestricted. Multiple individuals with food vendor licenses can legally operate from one permitted food vending unit. Merchandise licenses are renewed annually, while food licenses and permits are renewed biennially.
@@
2.2 Survey Design
The Street Vendor Project (SVP), part of the Urban Justice Center, is a non-profit organization that advocates on behalf of New York City street vendors. SVP administered a survey to approximately 2,000 street vendors while distributing Coronavirus relief aid in 2021. The aid was a one-time payment of \$1,000, available to any individual that owned, operated, or was otherwise employed by a street vending business in New York City between 2020 and 2021. First Amendment vendors were also eligible. There were no limits based on the size of the business, the number of sales, or whether licensed and/or permitted. All individuals eligible for aid were invited to complete a survey.
@@ 2.2 Survey Design
Of the 1,400 food vendors, 349 (25%) indicated they had a permit to vend. Of the 559 merchandise vendors, 505 were not veterans, of which 308 (61%) indicated they had a license to vend. The number of respondents is listed by neighborhood in Table 1. (Neighborhoods with few respondents are grouped together.) A map of the number of respondents by ZIP Code Tabulation Area (ZIP Code) is shown in the top left panel of Figure 1. Approximately five percent of vendors did not identify their vending location. These vendors are included in the New York City total, but they are excluded from the neighborhood estimates in Table 1 and Figure 1. (For this reason, the Respondent and Population columns do not sum exactly to the New York City total.) Veteran merchandise vendors are excluded entirely from Table 1 and Figure 1, although these vendors are reflected in our overall estimate of 23,000 vendors.
@@ 3. Methodology
We estimate the number of street vendors that operate in New York City. We use ratio estimation, which leverages the fact that the number of licenses and permits are limited by local law. In Section 3.1, we provide a simple explanation of ratio estimation, which is intended to be accessible to a general readership and highlights our main assumption. In Section 3.2, we develop a theoretical justification based on the theory of point processes, which, while more technical, allows us to explicitly state the estimands and underlying assumptions and to derive the standard errors. Inference is conducted using increasing domain asymptotics, and the details are outlined in Sections 7.1, 7.2, and 7.3.
@@ 3.1 Ratio Estimation
Ratio estimation is a common approach for estimating the size of a population. See Cochran (1978) and Hald (1998, chap. 16) for a historical discussion and Lohr (2021, chap. 4) for an introduction. We provide a simple explanation based on cross-multiplication, also called the rule of three. The purpose is to highlight the main assumption. A more rigorous derivation is presented in Section 3.2.
@@
$\hat{\Lambda}_{0}(A)=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}$
It is called the ratio estimator because it depends on the ratio of random variables $n_{0}(A)\,/\,n_{1}(A)$. (The total number of permitted vendors, $N_{1}(A)$, is considered fixed.) The estimator is approximately equal to $N_{0}(A)$ when the response rate, $p$, is approximately the same for each status, i.e.,
$p\approx\frac{n_{0}(A)}{N_{0}(A)}\approx\frac{n_{1}(A)}{N_{1}(A)}$
in which case $\hat{\Lambda}_{0}(A)$ is the solution for $N_{0}(A)$ in terms of $N_{1}(A)$, $n_{1}(A)$, and $n_{0}(A)$, obtained by cross-multiplication. It follows that $\hat{\tau}(A)=\hat{\Lambda}_{0}(A)+N_{1}(A)$ is approximately equal to the total $N_{0}(A)+N_{1}(A)$.
The main assumption of ratio estimation is that the survey data are representative in the sense that the response rate does not depend on whether a mobile food vendor has a permit. If we further assume that the response rate does not vary by subregion, we can also determine the total number of mobile food vendors by subregion. For example, suppose the response rate in subregion $B\subset A$ is also approximately equal to $p$, i.e.,
$p\approx\frac{n_{1}(A)}{N_{1}(A)}\approx\frac{n_{0}(B)}{N_{0}(B)}$
@@
$\tilde{\Lambda}_{0}(B)=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}$
@@
The subregion estimator is useful when $N_{1}(B)$ is not observed and $\hat{\Lambda}_{0}(B)$ cannot be calculated directly. This is the case with the survey data described in Section 2.2. We observe $N_{1}(A)$, the number of permits city-wide, but not $N_{1}(B)$, the number of permits within subregion $B$. (Recall that the vending locations of the respondents are recorded and thus $n_{0}(B)$ and $n_{1}(B)$ are observed for any $B\subset A$.)
@@
$p\approx\frac{n_{1}(A)}{N_{1}(A)}\approx\frac{n_{1}(B)}{N_{1}(B)}$
This suggests the estimator
@@ 3.2 Model
The Poisson process assumption may be justified by the law of rare events. Any sufficiently large area can be partitioned into a large number of theoretically vendable locations. Whether a vendor is located within a partition has a vanishingly small probability, such that the number of vendors in that area is well approximated by a Poisson distribution. This approximation is accurate even under weak dependence (Freedman 1974). Other justifications are possible. For example, see Section 4.2 for a justification of the Poisson process based on the theory of birth processes.
@@
[…]
The data are the locations of the respondents by status (i.e., the realizations of $\pi_{i}$) and the number of vendors in region $A$ with a permit, $N_{1}(A)$. In other words, we assume $n_{0}(B)$ and $n_{1}(B)$ are observed for any $B\subseteq A$, while $N_{0}(B)$ and $N_{1}(B)$ are unobserved with the exception of $N_{1}(A)$. Our goal is to estimate the expected number of street vendors in subregion $B\subseteq A$ without a permit,
[…]
and the expected total with or without a permit,
$\tau(B)=\Lambda_{0}(B)+\mathbb{E}\left[N_{1}(B)\,|\,N_{1}(A)\right]$
We condition on $N_{1}(A)$ to reflect the fact that while the number of vendors of either status in any subregion $B\subsetneq A$ is random, the total number of vendors with permits citywide is fixed by law.
In Section 3.1, we described the main assumption of ratio estimation: that the status and spatial distribution of the respondents is representative. We now make this statement precise. We assume that $p_{i}(x)$ can be decomposed into a constant that does not depend on $i$ or $x$ plus a spatially varying error term, $p_{i}(x)=p\,+\,\epsilon_{i}(x)$, and the error term is orthogonal to the corresponding intensity. i.e.,
[…]
It follows that
[…]
$n_{0}(B)\sim$ $\ \text{Poisson}\big{(}p\,\Lambda_{0}(B)\big{)}$
$n_{1}(B)\,|\,N_{1}(A)\sim$ $\ \text{Binomial}\big{(}N_{1}(B),\,p\,q(B)\big{)}$
where
$q(B)=\frac{\Lambda_{1}(B)}{\Lambda_{1}(A)}$
The maximum likelihood estimates for $\Lambda_{0}(A)$ and $p$ are
$\hat{\Lambda}_{0}(A)=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}\qquad\text{and}\qquad\hat{p}=\dfrac{n_{1}(A)}{N_{1}(A)}$
Using increasing domain asymptotics, $\hat{\Lambda}_{0}(A)$ is asymptotically normal with mean $\Lambda_{0}(A)$ and standard error
$\hat{\text{SE}}[\hat{\Lambda}_{0}(A)]=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(A)}+\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}}$
The maximum likelihood estimate for the total $\tau(A)$ is $\hat{\tau}(A)=\hat{\Lambda}_{0}(A)+N_{1}(A)$, which has the same standard error as $\hat{\Lambda}_{0}(A)$. See Section 7.1.2 for details.
The maximum likelihood estimate for $\Lambda_{0}(B)$ is
$\tilde{\Lambda}_{0}(B)=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}$
$\tilde{\text{SE}}[\tilde{\Lambda}_{0}(B)]=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(B)}+\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}}$
$\tilde{\tau}(B)=\dfrac{N_{1}(A)\,\left(n_{0}(B)+n_{1}(B)\right)}{n_{1}(A)}$
The asymptotic standard error is
$\tilde{\text{SE}}[\tilde{\tau}(B)]=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(B)}+\frac{1}{n_{1}(A)}-\frac{1}{N_{1}(A)}+\dfrac{n_{1}(B)\,n_{1}(A\backslash B)}{n_{1}(A)\,n_{0}(B)^{2}}}$
where $A\backslash B$ denotes the complement of subregion $B$. See Section 7.1.3 for details.

### v6

# Estimating the Number of Street Vendors in New York City:

@@ Abstract
We estimate the number of street vendors in New York City. First, we summarize the process by which vendors receive licenses and permits to operate legally in New York City. We then describe a survey that was administered by the Street Vendor Project while distributing coronavirus relief aid to vendors operating in New York City both with and without a license or permit. Finally, we review ratio estimation and develop a theoretical justification based on the theory of point processes. We find approximately 23,000 street vendors operate in New York City—20,500 mobile food vendors and 2,400 general merchandise vendors—with one third located in just six ZIP Codes—11368 (16%), 11372 (3%), and 11354 (3%) in North and West Queens and 10036 (5%), 10019 (4%), and 10001 (3%) in the Chelsea and Clinton neighborhoods of Manhattan. Our estimates suggest the American Community Survey misses the majority of New York City street vendors.
@@ 1. Introduction
Street vendors are New York City’s smallest businesses, selling food and merchandise from carts, stalls, and trucks throughout the five boroughs. They are an iconic part of the urban landscape. They are also a thriving sector of the local economy, contributing millions of dollars in government revenue through taxes, fines, and fees. Perhaps most importantly, street vending historically benefits underserved communities, both because vendors operate in neighborhoods with limited access to traditional stores and because vending is one of a handful of occupations in which New Yorkers of all backgrounds, immigrants in particular, can achieve economic mobility and a chance at the American dream (Burrows and Wallace 1998, chap. 42).
@@
In Section 3, we use the fact that the number of licenses and permits are limited by law to construct a ratio estimator for the number of street vendors. We then develop a theoretical justification based on the theory of point processes, which assumes the spatial distribution of survey respondents is representative and well approximated by a set of inhomogeneous Poisson processes. Finally, we derive the standard errors.
In Section 4, we examine the sensitivity of the ratio estimator to deviations from the underlying assumptions. We first consider the scenario in which the spatial distribution of the survey respondents is not representative, and we examine a more general estimator in which sampling weights adjust for differential nonresponse. We then consider the scenario in which the Poisson process assumption is violated, and the negative binomial process provides a better approximation. For both scenarios, we rederive the standard errors.
In Section 5, we estimate the number of street vendors in New York City. We find approximately 23,000 street vendors operate in New York City: 20,500 mobile food vendors and 2,400 general merchandise vendors. We also find that this estimate is somewhat sensitive to deviations from our underlying assumptions. For example, we find that if the spatial distribution of survey respondents is not representative, the estimated number of street vendors in New York City could be as low as 17,200 and as high as 41,800.
In Section 6, we conclude with a discussion. We first consider how the theoretical arguments provided in Sections 3 and 4 could be extended to more complex models. We then consider some implications of the estimates provided in Section 5. We find that although the spatial distribution of respondents is consistent with the results of the American Community Survey, the two disagree on the size of New York City’s street vending population. Our estimates suggest the American Community Survey misses the majority of New York City street vendors. We believe these findings shed light on the broader challenges of measuring the “gig economy.”
2. Background and Survey Operations
A street vendor is any individual who sells goods from a mobile vending unit instead of a store. We distinguish between the vending establishment—the mobile vending unit from which the goods are sold—and the vendors—the individuals who own and/or operate that unit. Note that an establishment refers to a single vending unit. There may be multiple vendors associated with any establishment, and multiple establishments may be owned by a single individual or firm. For example, suppose Sally owns the business Sally’s Salads, which consists of two food trucks, each staffed by three employees. Then Sally’s Salads counts as one business, two establishments, and seven vendors including Sally.
@@ 2.1 Background
A long list of rules determines the individuals, locations, and time periods during which vendors can legally sell their goods in New York City. Most relevant is the fact that a merchandise vendor requires a license to operate legally, and a food vendor requires both a license and a permit. The number of food permits and merchandise licenses are limited by law: 5,100 permits and 853 general merchandise licenses are available at the time of our writing. Food licenses are not limited. Note that a food permit is issued to an individual or business to allow for the sale of food from a specific mobile food vending unit, such as a cart or truck. A food license authorizes an individual to prepare or serve food from a permitted mobile food vending unit.
There are several versions of a food permit with varying restrictions. Of the 5,100 permits available, 200 are borough permits that limit vendors to one of the five boroughs; 100 are reserved for veterans or vendors with a disability; 1,000 are seasonal and valid only from April to October; 1,000 are green cart permits that limit vendors to selling fruit, vegetables, plain nuts, and water; and 2,800 are unrestricted. Multiple individuals with food licenses can legally operate from one permitted food vending unit. Merchandise licenses are renewed annually, while food licenses and permits are renewed biennially.
@@
2.2 Survey Operations
The Street Vendor Project (SVP), part of the Urban Justice Center, is a non-profit organization that advocates on behalf of New York City street vendors. SVP administered a survey to approximately 2,000 street vendors while distributing coronavirus relief aid in 2021. The aid was a one-time payment of \$1,000, available to any individual that owned, operated, or was otherwise employed by a street vending business in New York City between 2020 and 2021. First Amendment vendors were also eligible. There were no restrictions based on residency, the size of the business, the number of sales, or whether licensed and/or permitted. All individuals eligible for aid were invited to complete a survey.
@@ 2.2 Survey Operations
Of the 1,400 food vendors, 349 (25%) indicated they had a permit to vend. Of the 559 merchandise vendors, 505 were not veterans, of which 308 (61%) indicated they had a license to vend. The number of respondents is listed by neighborhood in Table 1. Neighborhoods with few respondents are grouped together. A map of the number of respondents by ZIP Code Tabulation Area (ZIP Code) is shown in the top left panel of Figure 1. Approximately five percent of vendors did not identify their vending location. These vendors are included in the New York City total, but they are excluded from the neighborhood estimates in Table 1 and Figure 1. For this reason, the Respondent and Population columns do not sum exactly to the New York City total. Veteran merchandise vendors are excluded entirely from Table 1 and Figure 1, although these vendors are reflected in our overall estimate of 23,000 vendors.
@@ 3. Methodology
We estimate the number of street vendors that operate in New York City. We use ratio estimation, which leverages the fact that the number of licenses and permits are limited by local law. In Section 3.1, we provide a simple explanation of ratio estimation. The explanation is intended to be accessible to a general readership and highlights our main assumption. In Section 3.2, we develop a theoretical justification based on the theory of point processes, which while more technical, allows us to explicitly state the estimands and underlying assumptions and to derive the standard errors. The details are outlined in Sections 7.1, 7.2, and 7.3.
@@ 3.1 Ratio Estimation
Ratio estimation is a common approach for estimating the size of a population. See Cochran (1978) and Hald (1998, chap. 16) for a historical discussion and Lohr (2021, chap. 4) for an introduction. We provide a simple explanation based on cross-multiplication, also called the rule of three. The purpose is to highlight the main assumption.
@@
$\hat{\Lambda}_{0}(A)=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}.$
It is called the ratio estimator because it depends on the ratio of random variables $n_{0}(A)\,/\,n_{1}(A)$. The total number of permitted vendors, $N_{1}(A)$, is considered fixed. We formally define the estimand that $\hat{\Lambda}_{0}(A)$ estimates in Section 3.2. For now, we note that the estimator is approximately equal to $N_{0}(A)$ when the response rate, $p$, is approximately the same for each status, i.e.,
$p\approx\frac{n_{0}(A)}{N_{0}(A)}\approx\frac{n_{1}(A)}{N_{1}(A)}.$
The formula for $\hat{\Lambda}_{0}(A)$ is obtained by using cross-multiplication to solve this equation for $N_{0}(A)$ in terms of $N_{1}(A)$, $n_{1}(A)$, and $n_{0}(A)$. It follows that $\hat{\tau}(A)=\hat{\Lambda}_{0}(A)+N_{1}(A)$ is approximately equal to the total $N_{0}(A)+N_{1}(A)$.
The main assumption of ratio estimation is that the survey data are representative in the sense that the response rate does not depend on whether a mobile food vendor has a permit. If we further assume that the response rate does not vary by subregion, we can also determine the total number of mobile food vendors by subregion. For example, suppose the response rate in subregion $B\subseteq A$ is also approximately equal to $p$, i.e.,
$p\approx\frac{n_{1}(A)}{N_{1}(A)}\approx\frac{n_{0}(B)}{N_{0}(B)}.$
@@
$\tilde{\Lambda}_{0}(B)=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)},$
@@
The subregion estimator is useful when $N_{1}(B)$ is not observed and $\hat{\Lambda}_{0}(B)$ cannot be calculated directly. This is the case with the survey data described in Section 2.2. We observe $N_{1}(A)$, the number of permits citywide, but not $N_{1}(B)$, the number of permits within subregion $B$. Recall that the vending locations of the respondents are recorded and thus $n_{0}(B)$ and $n_{1}(B)$ are observed for any $B\subseteq A$.
@@
$p\approx\frac{n_{1}(A)}{N_{1}(A)}\approx\frac{n_{1}(B)}{N_{1}(B)}.$
This suggests the subtotal estimator
@@ 3.2 Model
The Poisson process assumption may be justified by the law of rare events. Any sufficiently large area can be partitioned into a large number of theoretically vendable locations. Whether a vendor is located within a partition has a vanishingly small probability such that the number of vendors in that area is well approximated by a Poisson distribution. This approximation is accurate even under weak dependence (Freedman 1974). Other justifications are possible. See Section 4.2 for a justification of the Poisson process based on the theory of birth processes.
Let $\lambda_{i}(x)$ denote the intensity of the process $\Pi_{i}$ at location $x$ so that the number of individuals in any subregion $B\subseteq A$, $N_{i}(B)$, is distributed Poisson with mean measure $\Lambda_{i}(B)=\int_{B}\,\lambda_{i}(x)\,dx$, i.e.,
$N_{i}(B)\sim\text{Poisson}\left(\int_{B}\,\lambda_{i}(x)\,dx\right).$
We assume each individual responds independently to the survey with probability $p_{i}(x)>0$ such that by Campbell’s Theorem (Kingman 1992), the respondents constitute a “thinned” inhomogeneous Poisson process, $\pi_{i}$, with intensity $\lambda_{i}(x)\,p_{i}(x)$ at location $x$. The number of respondents, $n_{i}(B)$, is then distributed Poisson with mean $\int_{B}\,\lambda_{i}(x)\,p_{i}(x)\,dx$, i.e.,
$n_{i}(B)\sim\text{Poisson}\left(\int_{B}\,\lambda_{i}(x)\,p_{i}(x)\,dx\right).$
The data are the locations of the respondents by status (i.e., the realizations of $\pi_{i}$) and the number of vendors in region $A$ with a permit, $N_{1}(A)$. In other words, we assume $n_{0}(B)$ and $n_{1}(B)$ are observed for any $B\subseteq A$, while $N_{0}(B)$ and $N_{1}(B)$ are unobserved with the exception of $N_{1}(A)$.
Our goal is to estimate $\mathbb{E}\left[N_{0}(B)\,|\,N_{1}(A)\right]$, the expected number of street vendors in subregion $B\subseteq A$ without a permit, as well as $\tau(B)=\mathbb{E}\left[N_{0}(B)+N_{1}(B)\,|\,N_{1}(A)\right]$, the expected total with or without a permit. We always condition on $N_{1}(A)$ to reflect the fact that while the number of vendors of either status in any subregion $B\subsetneq A$ is random, the total number of vendors with permits citywide is fixed by law. However, we sometimes suppress the fact that $N_{1}(A)$ is fixed in our notation. For example, $N_{0}(B)$ is assumed to be independent of $N_{1}(A)$ so we often write $\mathbb{E}\left[N_{0}(B)\,|\,N_{1}(A)\right]=\Lambda_{0}(B)$ and $\tau(B)=\Lambda_{0}(B)+\mathbb{E}\left[N_{1}(B)\,|\,N_{1}(A)\right]$.
In Section 3.1, we described the main assumption of ratio estimation: the status and spatial distribution of the respondents is representative. We now make this statement precise. We assume that $p_{i}(x)$ can be decomposed into a constant that does not depend on $i$ or $x$ plus a spatially varying error term, $p_{i}(x)=p\,+\,\epsilon_{i}(x)$ and that the error term is orthogonal to the corresponding intensity. i.e.,
$\langle\lambda_{i},\epsilon_{i}\rangle=\int_{B}\lambda_{i}(x)\,\epsilon_{i}(x)\,dx=0.$
It follows that
$n_{i}(B)\sim\text{Poisson}\left(p\int_{B}\lambda_{i}(x)\,dx\right)$
$n_{0}(B)\sim$ $\ \text{Poisson}\big(p\,\Lambda_{0}(B)\big)$
$n_{1}(B)\,|\,N_{1}(A)\sim$ $\ \text{Binomial}\big(N_{1}(A),\,p\,q(B)\big)$
where
$q(B)=\frac{\Lambda_{1}(B)}{\Lambda_{1}(A)}.$
The maximum likelihood estimates for $\Lambda_{0}(A)$ and $p$ are
$\hat{\Lambda}_{0}(A)=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}\qquad\text{and}\qquad\hat{p}=\dfrac{n_{1}(A)}{N_{1}(A)}.$
The estimator $\hat{\Lambda}_{0}(A)$ is asymptotically normal with mean $\Lambda_{0}(A)$ and standard error
$\hat{\text{SE}}[\hat{\Lambda}_{0}(A)]=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(A)}+\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}}.$
The maximum likelihood estimate for the total $\tau(A)$ is $\hat{\tau}(A)=\hat{\Lambda}_{0}(A)+N_{1}(A)$, which is also asymptotically normal with mean $\tau(A)$ and the same standard error as $\hat{\Lambda}_{0}(A)$. See Section 7.2 for details.
The maximum likelihood estimate for $\Lambda_{0}(B)$ is
$\tilde{\Lambda}_{0}(B)=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)},$
$\tilde{\text{SE}}[\tilde{\Lambda}_{0}(B)]=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(B)}+\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}}.$
$\tilde{\tau}(B)=\dfrac{N_{1}(A)\,\left(n_{0}(B)+n_{1}(B)\right)}{n_{1}(A)},$
which is also asymptotically normal with mean $\tau(B)$ and standard error
@@
Note that while we derive asymptotic results for the expected number of vendors, a prediction interval for $N_{0}(B)$ or $N_{0}(B)+N_{1}(B)$ given $N_{1}(A)$ can also be constructed by noting that the distribution of $N_{i}(B)$ is approximately Poisson with rate $\hat{\Lambda}_{i}(B)$.